OpenStudy (anonymous):

The equation of the line passing through the point (4,2) with slope -3 is A. 3x+y-14=0 B. 3x+y+10=0 C. 3x+y-10=0 D. 3x+y+14=0 How do you find the answer?

OpenStudy (anonymous):

use y=mx+c then y=-3x+c put x=4 and y=2 2=-3*4+c c=2+12=14 y=2x+14 is your equation

OpenStudy (anonymous):

i know that slope is m=yB-yA/xB-xA

OpenStudy (anonymous):

but I don't know whats the point slope

OpenStudy (anonymous):

those are not needed in this but its good to know them the one you need to know is slope intercept formula

OpenStudy (anonymous):

okay slope formula y-y1 ----- =m x-x1 point slope formula y-y1 = m ( x - x1) slope intercept formula y= mx + b

OpenStudy (anonymous):

lets assume y = mx +c here we have m = slope = -3 so y = -3x+c lets put the value of the point in this to get the value of c 2= -3*4+c =>c=14 hence y = -3x+14 3x+y-14 a

OpenStudy (anonymous):

why is =>c=14?

OpenStudy (anonymous):

slope formula is y=mx+b this is what everyone is familar with ive never heard of y=mx+c y=-3x+b plug in the points and find ur answer

OpenStudy (anonymous):

okay thanks! :)

OpenStudy (anonymous):

No problem :)